Calculate the average acceleration needed by a train travelling initially at 12m/s[E] to stop in 120s.


1
Expert's answer
2022-03-08T08:55:50-0500

By definition, the acceleration is given as follows:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 0m\/s" is the final speed, "v_i = 12m\/s" is the initial speed, and "t = 120s" is the time.

Thus, obtain:


"a = \dfrac{0m/s-12m/s}{120s} = -0.1m/s^2"

If [E] is positive direction, then "a" is directed [W] because of "-" sign.


Answer. "a = 0.1m\/s^2[W]".